Test the hypothesis using the​ P-value approach.
Upper H 0 : p equals 0.50 versus Upper H 1 : p less than 0.50
n equals 150 comma x equals 66 comma alpha equals 0.10
Perform the test using the​ P-value approach.
whats that p value

Answer by Boreal(15235)

p hat is 66/150=0.44
z=(p hat-p)/sqrt (0.5*0.5/150)
=-0.06/0.0408=-1.47
p-value for that is 0.14